father age is 4 times that of his son. 5yrs back, it was 7 times. his age now is?
Answer by Cromlix(4381)

father age is 4 times that of his son.
F = 4S
5yrs back, it was 7 times.
F - 5 = 7(S - 5)
Replace F with 4S from first equation
4S - 5 = 7(S - 5)
4S - 5 = 7S - 35
Collect like terms
4S - 7S = -35 + 5
-3S = -30
S = 10.
Father's age = 40 years old
Son's age = 10 years old.
Hope this helps.
;-)